### v2

# thanks: Corresponding author

Abstract
Using the recently proposed covariant framework of general relativistic stochastic mechanics and stochastic thermodynamics, we proved the detailed and integral fluctuation theorems in curved spacetime. The time-reversal transformation is described as a transformation from the perspective of future-directed observer to that of the corresponding past-directed observer, which enables us to maintain general covariance throughout the construction. The result presented in this work may be applied in understanding the origin of irreversibility in macroscopic processes in the presence of relativistic gravity, as in most astrophysical or cosmological processes.
Langevin equation, fluctuation theorem, time-reversal symmetry, general relativity
Introduction. – A fundamental problem in modern statistical physics is the emergence of macroscopic irreversibility in systems which have time-reversal symmetry (TRS) in the underlying microscopic description. The relevant researches can be traced back to Boltzmann, who employed a scattering model with TRS in deriving the H-theorem [1], which states that the entropy of a macroscopic system cannot decrease in the course of time. However, the H-theorem has been under debates and challenges ever since its birth. The most famous challenge is known as Loschmidt paradox [2, 3], which argues that, if an entropy-increasing process exists for a system, the underlying TRS should also permit a corresponding entropy-decreasing process. The fluctuation theorems, proposed in the 1990s [4, 5, 6], offer an interpretation for Loschmidt paradox. Basically, these theorems state that, while both entropy-increasing and entropy-decreasing processes are permissible, the former is more probable.
In the context of stochastic mechanics, Sekimoto [7] utilized the overdamped Langevin equation to properly classify the energy exchange of a Brownian particle with the heat reservoir into trajectory heat and work, and thus establishes the first law of stochastic thermodynamics on the trajectory level. This allows for the construction of fluctuation theorems based on stochastic mechanics [8, 9, 10, 11, 12].
The initial studies on fluctuation theorems are mostly carried out in non-relativistic theories. Since 2007, some attempts [13, 14, 15, 16, 17, 18] in establishing fluctuation theorems in the special relativistic context appeared. Such attempts are important because relativity imposes stronger protection on the spacetime symmetry, making it harder to break the TRS. The purpose of the present work is to broaden the scope of the fluctuation theorem to encompass curved spacetime while maintaining general covariance. This is also important, because, on the one hand, gravity is a universal interaction, it is desirable to see whether gravity has any impact on the origin of irreversibility — another universal phenomenon that appear in the scope of macroscopic theories; on the other hand, the choice of time parameter is more subtle in general relativistic theories than in special relativistic cases.
Conventionally, the time-reversal transformation (TRT) is merely described as a transformation of the time parameter, $t\mapsto-t$. However, if the time parameter is identified to be the coordinate time $x^{0}$, such a transformation will result in a lack of covariance. Our covariant framework [19, 20] for relativistic stochastic mechanics urges that the choice of time parameter should be closely linked with the observer. Consequently, the TRT needs be realized as a change from the perspective of a future-directed observer to that of a past-directed observer. In such a realization, the coordinate system is left intact.
Relativistic Langevin dynamics. – Langevin equation describes the motion of a heavy particle, referred to as the Brownian particle, under the random disturbance of a heat reservoir. Conventionally, Langevin equation can be expressed in the form of Newton’s second law, incorporating elements such as the random force, damping force, and various other external forces. For a relativistic particle of mass $m$ and charge $q$ moving in $(d+1)$-dimensional spacetime $\mathcal{M}$ with the metric $g_{\mu\nu}$, the Langevin equation employing the particle’s proper time $\tau$ as evolution parameter is is referred to as LEτ [19],
${\rm d}\tilde{x}_{\tau}^{\mu}$ $=\frac{\tilde{p}_{\tau}^{\mu}}{m}{\rm d}\tau,$ (1)
${\rm d}\tilde{p}_{\tau}^{\mu}$ $=\xi^{\mu}_{\tau}{\rm d}\tau+\mathcal{F}_{\text{dp}}^{\mu}{\rm d}\tau-\frac{1}{m}\varGamma^{\mu}{}_{\alpha\beta}\tilde{p}_{\tau}^{\alpha}\tilde{p}_{\tau}^{\beta}{\rm d}\tau+\mathcal{F}_{\text{em}}^{\mu}{\rm d}\tau,$ (2)
where $\xi^{\mu}_{\tau}$ is the random force, $\mathcal{F}_{\text{dp}}^{\mu}:=\mathcal{K}^{\mu\nu}U_{\nu}$ is the damping force in which the damping coefficient tensor $\mathcal{K}^{\mu\nu}$ obeys $\mathcal{K}^{\mu}{}_{\nu}p^{\nu}=0$, $\mathcal{F}_{\text{em}}^{\mu}:=\frac{q}{m}F^{\mu}{}_{\nu}\tilde{p}^{\nu}_{\tau}$ is the electromagnetic force, and $\varGamma^{\mu}{}_{\alpha\beta}$ denotes the Christoffel connection associated with $g_{\mu\nu}$. Tilded symbols such as $\tilde{x}_{\tau}$ and $\tilde{p}_{\tau}$ represent random variables, and the un-tilded ones represent their realization. The random force is consisted of a term encoding the Stratonovich coupling between the stochastic amplitudes $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ (which transforms as a vector field for each fixed $\mathfrak{a}$) with a set of $d$ independent Wiener process ${\rm d}\tilde{w}_{\tau}^{\mathfrak{a}}$ obeying the probability distribution
$\Pr[{\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}={\rm d}w^{\mathfrak{a}}]=\frac{1}{(2\pi{\rm d}\tau)^{d/2}}\exp\left[-\frac{\delta_{\mathfrak{a}\mathfrak{b}}{\rm d}w^{\mathfrak{a}}{\rm d}w^{\mathfrak{b}}}{2{\rm d}\tau}\right]$ (3)
of variance ${\rm d}\tau$, together with a term incorporating an additional stochastic force $\mathcal{F}^{\mu}_{\text{add}}:=\frac{\delta^{\mathfrak{a}\mathfrak{b}}}{2}\mathcal{R}^{\mu}{}_{\mathfrak{a}}\nabla^{(h)}_{i}\mathcal{R}^{i}{}_{\mathfrak{b}}$,
$\xi_{\tau}^{\mu}:=\mathcal{R}^{\mu}{}_{\mathfrak{a}}\circ_{S}{\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}/{\rm d}\tau+\mathcal{F}^{\mu}_{\text{add}},$ (4)
where $\nabla^{(h)}$ is the covariant derivative on the mass shell and $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ stands for the stochastic amplitude which may depend on $\tilde{x}_{\tau}^{\mu}$ and $\tilde{p}_{\tau}^{\mu}$ and transform as a vector field for each fixed $\mathfrak{a}$. The Stratonovich coupling $\circ_{S}$ maintains the chain rule in stochastic calculus, making the general covariance of LEτ self-evident.
The energy of a relativistic particle measured by the observer with proper velocity $Z^{\mu}$ is defined as
$E(x,p):=-Z_{\mu}(x)p^{\mu}.$ (5)
As the microstate of the Brownian particle evolves in a stochastic process, the energy of the Brownian particle also becomes a stochastic process $\tilde{E}_{\tau}:=E(\tilde{x}_{\tau},\tilde{p}_{\tau})$. The chain rule implies that
${\rm d}\tilde{E}_{\tau}$ $=\frac{\partial E}{\partial x^{\mu}}{\rm d}\tilde{x}^{\mu}_{\tau}+\frac{\partial E}{\partial p^{\mu}}{\rm d}\tilde{p}^{\mu}_{\tau}$
$=-Z_{\mu}\left[\xi_{\tau}^{\mu}+\mathcal{F}_{\text{dp}}^{\mu}\right]{\rm d}\tau-\frac{\tilde{p}_{\tau}^{\mu}\tilde{p}_{\tau}^{\nu}}{m}\nabla_{\nu}Z_{\mu}{\rm d}\tau-Z_{\mu}\mathcal{F}_{\text{em}}^{\mu}{\rm d}\tau.$ (6)
The second and last terms in eq. (General Relativistic Fluctuation Theorems) are recognized to be the gravitational [21] and electromagnetic works,
${\rm d}\tilde{\mathcal{P}}_{\tau}:=-\frac{\tilde{p}_{\tau}^{\mu}\tilde{p}_{\tau}^{\nu}}{m}\nabla_{\nu}Z_{\mu}{\rm d}\tau,\quad{\rm d}\tilde{\mathcal{W}}_{\tau}:=-Z_{\mu}\mathcal{F}^{\mu}_{\text{em}}{\rm d}\tau,$ (7)
and the first term incorporates the energy change caused by the impact of the heat reservoir, which is by definition the heat received by the Brownian particle,
${\rm d}\tilde{Q}_{\tau}:=-Z_{\mu}\left[\xi_{\tau}^{\mu}+\mathcal{F}_{\text{dp}}^{\mu}\right]{\rm d}\tau.$ (8)
Therefore, eq. (General Relativistic Fluctuation Theorems) becomes precisely the first law of general relativistic stochastic thermodynamics,
${\rm d}\tilde{E}_{\tau}={\rm d}\tilde{\mathcal{Q}}_{\tau}+{\rm d}\tilde{\mathcal{P}}_{\tau}+{\rm d}\tilde{\mathcal{W}}_{\tau}.$ (9)
Time-reversal symmetry. – Given a worldline $x_{\tau}$ for a massive particle in the spacetime $\mathcal{M}$, its tangent vector $p^{\mu}_{\tau}$ can be either aligned with the proper velocity $Z_{\mu}$ of some chosen observer Alice, i.e. $p^{\mu}_{\tau}Z_{\mu}<0$, or opposite to the proper velocity $C_{\mu}$ of some other observer, Carol, i.e. $p^{\mu}_{\tau}C_{\mu}>0$. We will refer to Alice as future-directed and to Carol as past-directed. To be more specific, we fix Carol to be the TRT image of Alice, i.e. $C_{\mu}=-Z_{\mu}$. If Alice perceives the worldline evolving from $t_{I}$ to $t_{F}$, then Carol will perceive the worldline evolving from $t_{F}$ to $t_{I}$. This leads to a change in the sign of the time derivatives, e.g. ${\rm d}/{\rm d}\tau\rightarrow-{\rm d}/{\rm d}\tau$, and consequently the momentum $p$ needs to reverse its sign. Notice however that ${\rm d}\tau$ resp. ${\rm d}t$ denote the size of infinitesimal temporal steps, both remain unchanged under the TRT.
The mass shell bundle $\Gamma_{m}:=\{(x,p)\in T\mathcal{M}|p_{\mu}p^{\mu}=-m^{2}\}$ can be separated into two non-path-connected regions, i.e. the future mass shell bundle
$\Gamma^{+}_{m}:=\{(x,p)\in\Gamma_{m}|Z_{\mu}p^{\mu}<0\},$ (10)
and the past mass shell bundle
$\Gamma^{-}_{m}:=\{(x,p)\in\Gamma_{m}|Z_{\mu}p^{\mu}>0\},$ (11)
both are defined relative to Alice. Clearly, there is a homeomorphism between $\Gamma^{+}_{m}$ and $\Gamma^{-}_{m}$:
$I:\Gamma^{+}_{m}\to\Gamma^{-}_{m},\qquad(x,p)\mapsto(x,-p).$ (12)
This homeomorphism is the mathematical realization of the TRT. The phase trajectory $Y_{t}=(y_{t},k_{t})$ of the Brownian particle is the uplift of the worldline into $\Gamma^{+}_{m}$, and is referred to as the forward trajectory, while the TRT image of the phase trajectory $Y^{-}_{t}=I(Y_{t_{I}+t_{F}-t})$ is the uplift of the worldline into $\Gamma^{-}_{m}$, and is known as the reversed trajectory.
Now consider a macroscopic system consisting of a great number of massive charged particles which can be classified into several species. The electromagnetic field is produced by the charges carried by the particles and the geometry of the $(d+1)$-dimensional spacetime is determined by the masses and charges within the system. The phase trajectories of these particles obey the following equations,
$\frac{{\rm d}x^{\mu}_{s}}{{\rm d}\tau_{s}}$ $=\frac{p^{\mu}_{s}}{m_{s}},$ (13)
$\frac{{\rm d}p^{\mu}_{s}}{{\rm d}\tau_{s}}$ $=\frac{q_{s}}{m_{s}}F^{\mu}{}_{\nu}p^{\nu}_{s}-\frac{1}{m_{s}}\varGamma^{\mu}{}_{\alpha\beta}p^{\alpha}_{s}p^{\beta}_{s},$ (14)
where the suffices $s$ indicate different species. Assuming that each species of particles obeys a distinct TRT invariant distribution $\Phi_{s}(x_{s},p_{s})$, the electric current and the energy-momentum tensor contributed by the particles can be written as
$J^{\mu}$ $=\sum_{s}\int\varpi_{s}q_{s}p^{\mu}_{s}\Phi_{s},$ (15)
$\qquad T^{\mu\nu}_{\text{pa}}$ $=\sum_{s}\int\varpi_{s}p^{\mu}_{s}p^{\nu}_{s}\Phi_{s},$ (16)
where $\varpi_{s}:=(\sqrt{g}/(p_{s})_{0}){\rm d}p^{1}_{s}\wedge...\wedge{\rm d}p^{d}_{s}$ is the invariant volume element on the mass shell of the $s$-species. The Maxwell equation and the energy-momentum tensor produced by the electromagnetic field are given by
$\nabla_{\nu}F^{\mu\nu}=J^{\mu},\quad T^{\mu\nu}_{\text{em}}=F^{\mu\rho}F^{\nu}{}_{\rho}-\frac{1}{4}g^{\mu\nu}F^{\rho\sigma}F_{\rho\sigma},$ (17)
and finally, the Einstein equation that determines the spacetime geometry reads
$R^{\mu\nu}-\frac{1}{2}g^{\mu\nu}R=8\pi G\left(T^{\mu\nu}_{\text{pa}}+T^{\mu\nu}_{\text{em}}\right).$ (18)
As we shift the perspective from Alice to Carol, there exist multiple equivalent conventions [22, 23] regarding the TRT of the electromagnetic field and the charge. We adopt Feynman’s convention [24]:
$q\mapsto-q,\quad F^{\mu\nu}\mapsto F^{\mu\nu}.$ (19)
It becomes straightforward to verify that all equations of motion for the above system, i.e. eqs. (13)-(18), are invariant under the TRT. Therefore, in the presence of TRS, our designation of Alice as future-directed observer has no microscopic meaning: one cannot distinguish Alice from Carol purely from the mechanical description.
Stochastic mechanics, in essence, serves as an effective theory for complicated mechanical systems within specific spatial and temporal scales [25, 26, 27]. Consider a single heavy particle in the above system, and assuming that the remaining particles constitute a heat reservoir which have already attained thermal equilibrium. Due to the presence of gravity, the equilibrium state is not a state with uniform temperature, but rather a state of a fluid with the temperature obeying [28]
$\nabla_{\mu}T_{\rm B}+{T_{\rm B}}U^{\nu}\nabla_{\nu}U_{\mu}=0,$ (20)
where $U^{\mu}$ denotes the proper velocity of the fluid element. The non-uniform distribution of the temperature described by eq. (20) is known as Tolman-Ehrenfest effect. Please be reminded that, in relativistic context, the temperature is observer-dependent. The particular temperature $T_{\rm B}$ appeared above is the one perceived by the observer comoving with the heat reservoir, Bob, as indicated by the suffix $\rm B$.
The electromagnetic interaction exerted on the heavy particle can be separated into two components: the coarse-grained averaging effects at larger spatial and temporal scales, and the stochastic approximations at smaller scales. The latter encompasses both the random force and damping force. As a result, eqs. (13)-(14) can be approximated by LEτ [19], rendering this heavy particle as the Brownian particle. It will be clear shortly that the above coarse-grained picture breaks the TRS.
Fluctuation theorem. – Selecting an integral curve of Alice, we can interpret its arc length as the proper time $t$ of Alice, which can be further extended as a scalar field on $\mathcal{M}$ [19]. The proper time of Alice and that of the Brownian particle are connected via the following relation:
${\rm d}t=\frac{p^{\mu}}{m}\partial_{\mu}t{\rm d}\tau=\gamma(x,p){\rm d}\tau,$ (21)
where $\gamma(x,p)$ is the local Lorentz factor arising from the relative motion between Alice and the particle. Since the microstate $(x,p)$ of the Brownian particle is random, the infinitesimal increment of $\tau$ also becomes random from Alice’s perspective. Consequently, for Alice, LEτ needs to be reparameterized using the deterministic parameter $t$ in place of the random parameter $\tau$, yielding a new form of Langevin equation known as LEt [19],
${\rm d}\tilde{y}_{t}^{\mu}$ $=\frac{\tilde{k}^{\mu}_{t}}{m}\gamma^{-1}{\rm d}t,$ (22)
${\rm d}\tilde{k}_{t}^{\mu}$ $=\hat{\xi}_{t}^{\mu}\gamma^{-1}{\rm d}t+\mathcal{F}^{\mu}_{\text{dp}}\gamma^{-1}{\rm d}t$
$+\mathcal{F}^{\mu}_{\text{em}}\gamma^{-1}{\rm d}t-\frac{1}{m}\varGamma^{\mu}{}_{\alpha\beta}\tilde{k}^{\alpha}_{t}\tilde{k}^{\beta}_{t}\gamma^{-1}{\rm d}t,$ (23)
where $\tilde{Y}_{t}=(\tilde{y}_{t},\tilde{k}_{t})=(\tilde{x}_{\tilde{\tau}_{t}},\tilde{p}_{\tilde{\tau}_{t}})$, and
$\hat{\xi}^{\mu}_{t}:=\gamma^{1/2}\mathcal{R}^{\mu}{}_{\mathfrak{a}}\circ_{S}{\rm d}\tilde{W}_{t}^{\mathfrak{a}}/{\rm d}t+\mathcal{F}^{\mu}_{\text{add}}-\frac{1}{2}\mathcal{D}^{\mu i}\gamma^{1/2}\nabla^{(h)}_{i}\gamma^{-1/2},$ (24)
in which ${\rm d}\tilde{W}_{t}^{\mathfrak{a}}$ represents a Wiener process of variance ${\rm d}t$ for each fixed $\mathfrak{a}$. It is evident that LEt is observer-dependent, but still manifestly general covariant. The stochastic process $\tilde{Y}_{[t]}$ reparameterized by Alice is referred to as the forward process, while the stochastic process $\tilde{Y}^{-}_{[t]}$ reparameterized by Carol is referred to as the reversed process. It is worth noticing that process and trajectory are different concepts, the latter is a concrete realization of the former. In particular, the initial state of the reversed process and final state of the forward process are related via
$\tilde{Y}^{-}_{I}=I(\tilde{Y}_{F}),$ (25)
and there need not be any relationship between other states from the forward and reversed processes.
Contrary to conventional mechanical equations, Langevin equation lacks the capacity to determine whether a trajectory qualifies as its solution. Instead, the best one can do is to determine the probability of a given trajectory. Correspondingly, the TRS breaking manifests as the difference in the probabilities of the forward and reversed trajectories,
$\Pr[\tilde{Y}_{[t]}=Y_{[t]}]\neq\Pr[\tilde{Y}_{[t]}^{-}=Y_{[t]}^{-}].$ (26)
In order to quantify to what extent the TRS is broken in the general relativistic Langevin dynamics described by LEt, let us turn to the construction of the fluctuation theorem. The stochastic process governed by LEt is a Markov process. Consequently, the trajectory probability can be expressed as the product of the initial and conditional probabilities:
$\Pr[\tilde{Y}_{[t]}=Y_{[t]}]=\Pr[\tilde{Y}_{[t]}=Y_{[t]}|\tilde{Y}_{I}=Y_{I}]\Pr[\tilde{Y}_{I}=Y_{I}],$ (27)
and the conditional probability can be further decomposed into the product of the transition probabilities for a sequence of intermediate steps [20]. The initial probability $\Pr[\tilde{Y}_{I}=Y_{I}]$ is related to the one particle distribution $\varphi(Y_{I})$ of the relativistic Brownian particle via
$\Pr[\tilde{Y}_{I}=Y_{I}]=\gamma\lambda^{-1}\varphi(Y_{I}),\quad\lambda:=|\partial_{\mu}t|,$ (28)
where $\varphi(Y_{I})$ obeys the reduced Fokker-Planck equation [20], but is not a probability distribution by itself. The entropy density of the Brownian particle should be defined as [29]
$S:=-\ln\varphi,$ (29)
the difference of which at the final and the initial states of a trajectory is defined to be the trajectory entropy production. Furthermore, the ratio of the conditional probabilities of the forward and reversed trajectories is given by [23]:
$\frac{\Pr[\tilde{Y}_{[t]}=Y_{[t]}|\tilde{Y}_{I}=Y_{I}]}{\Pr[\tilde{Y}_{[t]}^{-}=Y^{-}_{[t]}|\tilde{Y}^{-}_{I}=Y^{-}_{I}]}$
$=$ $\frac{(\lambda\gamma^{-1})|_{Y_{I}}}{(\lambda\gamma^{-1})|_{Y_{F}}}\exp\left[\int^{t_{F}}_{t_{I}}\gamma^{-1}{\rm d}t\frac{1}{T_{\rm B}}(ma^{\mu}-\mathcal{F}^{\mu}_{\text{em}})U_{\mu}\right],$ (30)
where $a^{\mu}=p^{\nu}\nabla_{\nu}p^{\mu}/m^{2}$ is the proper acceleration of the Brownian particle. From the perspective of Bob, the complete differential of the energy is:
${\rm d}E=-\frac{p^{\nu}}{m}\nabla_{\nu}(p^{\mu}U_{\mu}){\rm d}\tau=-ma^{\mu}U_{\mu}{\rm d}\tau+{\rm d}\mathcal{P}.$ (31)
Comparing with eq. (9), it can be deduced that the integral over the exponent in eq. (General Relativistic Fluctuation Theorems) is equivalent to the entropy increase of the heat reservoir, thanks to the relativistic Clausius equality at the trajectory level:
$\int\gamma^{-1}{\rm d}t\frac{1}{T_{\rm B}}(ma^{\mu}-\mathcal{F}^{\mu}_{\text{em}})U_{\mu}=-\int\frac{{\rm d}\mathcal{Q}}{T_{\rm B}}=\Delta S_{R}.$ (32)
The relativistic Clausius equality at the ensemble level has been discussed in our previous work [29]. Although the relativistic Clausius equalities are only valid from the perspective of Bob, the amount of the entropy increase is observer-independent. Therefore, the ratio of the probabilities of the forward and reversed trajectories can be deduced from eq. (29) and eq. (General Relativistic Fluctuation Theorems) ,
$\frac{\Pr[\tilde{Y}_{[t]}=Y_{[t]}]}{\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]}=$ $\frac{\Pr[\tilde{Y}_{[t]}=Y_{[t]}|\tilde{Y}_{I}=Y_{I}]}{\Pr[\tilde{Y}_{[t]}^{-}=Y_{[t]}^{-}|\tilde{Y}_{I}^{-}=Y_{I}^{-}]}\frac{\Pr[\tilde{Y}_{I}=Y_{I}]}{\Pr[\tilde{Y}_{I}^{-}=Y_{I}^{-}]}$
$=$ ${\rm e}^{\Delta S_{R}}\frac{(\lambda\gamma^{-1})|_{Y_{I}}}{(\lambda\gamma^{-1})|_{Y_{F}}}\frac{\Pr[\tilde{Y}_{I}=Y_{I}]}{\Pr[\tilde{Y}_{F}=Y_{F}]}$
$=$ ${\rm e}^{\Delta S_{R}+\Delta S}.$ (33)
Let $\Sigma_{Y_{[t]}}=\Delta S_{R}+\Delta S$ be the total entropy production contributed by the trajectory $Y_{[t]}$, eq. (General Relativistic Fluctuation Theorems) can be rewritten as
$\frac{\Pr[\tilde{Y}_{[t]}=Y_{[t]}]}{\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]}={\rm e}^{\Sigma_{Y_{[t]}}}.$ (34)
This is the standard form of fluctuation theorem on the trajectory level. Furthermore, by integration over the space of trajectories and employing Jensen’s inequality, we arrive at the integral fluctuation theorem:
${\rm e}^{-\left\langle\Sigma_{\tilde{Y}_{[t]}}\right\rangle}$ $\leq\left\langle{\rm e}^{-\Sigma_{\tilde{Y}_{[t]}}}\right\rangle$
$=\int\mathcal{D}[Y_{[t]}]{\rm e}^{-\Sigma_{\tilde{Y}_{[t]}}}\Pr[\tilde{Y}_{[t]}=Y_{[t]}]$
$=\int\mathcal{D}[Y_{[t]}]\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]$
$=\int\mathcal{D}[Y^{-}_{[t]}]\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]=1,$ (35)
which means the statistical expectation value of the total entropy production must be no-negative, i.e. $\left\langle\Sigma_{\tilde{Y}_{[t]}}\right\rangle\geq 0$. For most trajectories, the total entropy production $\Sigma_{Y_{[t]}}$ is nonzero, indicating distinct probabilities for forward and reversed trajectories. The fluctuation theorem thus provides a quantitative link between TRS breaking and entropy production.
Conclusion. – Based on the covariant framework of general relativistic stochastic mechanics, we formulated the first law of stochastic thermodynamics in curved spacetime. By properly addressing the description for the TRS and its breaking, the corresponding fluctuation theorem is also proved in both the differential and integral forms. The observers and their behaviors under the TRT play a key role in our construction. The fact that the final fluctuation theorems (34) and (General Relativistic Fluctuation Theorems) take the same form as in non-relativistic context [8] is expected, because both the trajectory probability and the total entropy production are neither coordinate dependent nor observer dependent. The version of the fluctuation theorems proved in this work may shed some light on understanding the origin of irreversibility in the presence of relativistic gravitational interaction, as in most cases of astrophysical or cosmological processes.
Acknowledhement – This work is supported by the National Natural Science Foundation of China under the grant No. 12275138.
Conflict of interests declaration: – The authors declare no known conflict of interests.